OpenStudy (anonymous):

x^3-4x^2+4x=0 Solve for x

OpenStudy (tkhunny):

Have you considered factoring?

OpenStudy (anonymous):

\[x(x^2-4x+4) = x(x-2)^2 so x = 0, 2\] 2 is a solution with a mutliplicity of 2 meaning that 2 is a solution twice

OpenStudy (anonymous):

so then x =2

OpenStudy (anonymous):

I don't know where to begin I need step by step instructions

OpenStudy (anonymous):

Will someone please help me

OpenStudy (anonymous):

Let me type it out for you

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so first you factor out an x like I did above so you get \[x(x^2−4x+4)\] then you factor by grouping in this case you get \[x(x-2)^2\] set each of those equal to 0 \[x=0\] & \[(x-2)^2=0\] solve \[x=0,2\] and 2 with a multiplicity of 2 meaning that 2 is a solution twice. Make sense?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

Glad to help! :)